OpenStudy (anonymous):

Solve 15^2x = 36. Round to the nearest ten-thousandth.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

A. 0.6616 B. 2.6466 C. 1.7509 D. 1.9091

OpenStudy (whpalmer4):

Is that \[15^{2x} = 36\]?

OpenStudy (whpalmer4):

Yes, it must be, because I get one of those answers :-) Do you remember the property of logs that says \[\log (x^a) = a\log(x)\]? You can use that to solve this problem by taking the log of both sides and using that property.

OpenStudy (whpalmer4):

That means \[2x \log 15 = \log 36\]Now just solve for \(x\) and put the calculator to work

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

x=270 ?

OpenStudy (anonymous):

Ik it's wrong please let me explain ...

OpenStudy (anonymous):

was I supposed to do 36*15 ?

OpenStudy (whpalmer4):

\[2x \log 15 = \log 36\]Divide both sides by \(\log 15\) to get\[2x = \frac{\log 36}{\log 15}\]Can you do the rest? The answer had better be less than 1, because if \(x =1\), then \(15^{2x} = 15^2 = 225\) which is much bigger than 36!

OpenStudy (anonymous):

I got 1.323283792

OpenStudy (whpalmer4):

in fact, you could work it out in your head close enough to get the right answer! hmm. that's 2x the right answer...did you remember to divide by that 2 in front of \(x\)?

OpenStudy (anonymous):

A. 0.6616

OpenStudy (anonymous):

Thanks again

OpenStudy (anonymous):

one more please i promise and i will leave you alone ?

OpenStudy (whpalmer4):

So, you could estimate this as follows: \[15^{2x} = 36\]\[15^x*15^x = 36\]\[15^x = \sqrt{36}\]\[15^x = 6\]\[15^{1/2} = \sqrt{15} \approx \sqrt{16} = 4\]so it has to be bigger than 1/2 and smaller than 1. You could try 2/3: \[15^{2/3} = \sqrt[3]{15^2}\]and because \[15^2 = 225\] and \[6^3 = 216\]that means that \[15^{2/3}\approx 6\]and the guess of \(x = 2/3\) is pretty close!

OpenStudy (whpalmer4):

or, if you're like me, and memorized a handful of logarithms: \[15^{2x} = 36\]\[x = \frac{\log 36}{2\log 15}\]but \[36=6^2 = 2^2*3^2\]so \[\log 36 = 2\log 2+2\log3 = 2*0.30103 + 2*0.477121 \approx 1.5563\]and \[\log 15 = \log(3*5) = \log3 + \log 5 \approx 0.477121 + 0.69897 \approx 1.17609\]so our answer is \[x = \frac{1.5563}{2*1.17609} \approx 0.661642\]just like you got, except without a calculator (except to do the division)

OpenStudy (whpalmer4):

what's the other problem?